OpenStudy (anonymous):

Simplify 9 over quantity of 3 minus 2 i

OpenStudy (alexandervonhumboldt2):

is your problem \[\frac{ 9 }{ 3-2i }?\]

OpenStudy (anonymous):

Yes

OpenStudy (alexandervonhumboldt2):

mutlpy by a congugate: conngugete of 3-2i=3+2i

OpenStudy (danjs):

you mean put it into the complex number form?

OpenStudy (alexandervonhumboldt2):

\[\frac{ 9 }{ 3-2i }*\frac{ 3+2i }{ 3+2i}=\frac{ 9*(3+2i) }{ 3^2-(2i)^2 }=simplify\]

OpenStudy (alexandervonhumboldt2):

do you need to rationolize the denominator? or what? yeah

OpenStudy (anonymous):

5/5 = 40/ 5? Is that correct

OpenStudy (danjs):

yeah i^2=-1 put that into the complex number form prolly a + b*i

OpenStudy (danjs):

(27+18i) / (9 - 4(-1)) 27/13 + (18/13) *i

OpenStudy (anonymous):

The options are Simplify 9 over quantity of 3 minus 2 i. quantity of negative 27 minus 18 i over 13 9 i over quantity of 2 plus 3 i quantity of 27 plus 18 i over 5 quantity of 27 plus 18 i over 13

OpenStudy (alexandervonhumboldt2):

hint: \[\frac{ 27 }{ 13 }+\frac{ 18 }{ 13 }i=\frac{ 27+18i }{ 13 }\]

OpenStudy (anonymous):

I'm trying to figure it out

OpenStudy (anonymous):

How did you get 13?